What is the molecular equationi, ionic equation and net ionic equation of the reaction between aqueous chlorine and sodium iodide?


1
Expert's answer
2020-10-28T08:42:52-0400

Cl2(aq)+2NaI(aq)=2NaCl(aq)+I2(s)

Cl2(aq)+2Na+(aq)+2I-(aq)=2Na+(aq)+2Cl-(aq)+I2(s)

Cl2(aq)+2I-(aq)=2Cl-(aq)+I2(s)